SPECTRAL STUDY OF AMIKACIN - DNA INTERACTION

\underline{LOREDANA VÎJAN}, ALINA RĂDUCU and MIRCEA BĂRBUCEANU, University of Piteşti, Faculty of Science, Department of Physical Chemistry, Târgu din Vale 1, Piteşti, Romania.

Amikacin is an aminoglycoside antibiotic. Aminoglycoside antibiotics are bactericidal agents that are comprised of two or more amino sugars joined in glycosodic linkage to a hexoze nucleus. They contain a unique polyamine/ carbohydrate structure and have attracted considerable attention because of their specific interaction with nucleic acids. The bactericidal effects stem from their ability to block protein synthesis by binding to the A-site on ribosomal RNA. In this spectral study, the equilibrium binding constants were determined using visible absorption spectra. The results are rationalized in terms of several literature methods: Wolfe, Schwarz and Scatchard. The results indicate a tight binding between the drug and DNA.